# Lightweight Transformer via Unrolling of Mixed Graph Algorithms for Traffic Forecast

### v1

@@ 3.2.2 $\ell_{1}$-Norm Variational Term
Unlike DGLR in (2), DGTV is not symmetric and has no obvious frequency interpretation. Nonetheless, we provide a two-channel filterbank interpretation in the sequel.
@@ 3.3 Optimization Formulation
where $\mu_{u},\mu_{d,2},\mu_{d,1}\in\mathbb{R}_{+}$ are weight parameters for the three regularization terms. Combination of $\ell_{2}$- and $\ell_{1}$-norm penalties—DGLR and DGTV for directed graph ${\mathcal{G}}^{d}$ in our case [^3] We can employ both $\ell_{2}$- and $\ell_{1}$-norm regularization terms for the undirected graph ${\mathcal{G}}^{u}$ as well. For simplicity, we focus on designing new regularization terms for directed graphs in this paper. —is called elastic net regularization in statistics [27] with demonstrable improved robustness. Further, employing both regularization terms means that, after algorithm unrolling, we can adapt an appropriate mixture by learning weights $\mu_{d,2}$ and $\mu_{d,1}$ per neural layer.
(4) is a convex minimization problem composed of convex smooth $\ell_{2}$-norm terms and one convex non-smooth $\ell_{1}$-norm term. We pursue a divide-and-conquer approach and solve (4) via an ADMM framework [21] by minimizing $\ell_{2}$- and $\ell_{1}$-norm terms in (4) alternately.
@@ 3.4 ADMM Optimization Algorithm
We first introduce auxiliary variable ${\boldsymbol{\phi}}$ and rewrite the optimization in (4) as
@@
Using the augmented Lagrangian method [28], we rewrite (5) in an unconstrained form:
$\min_{{\mathbf{x}},{\boldsymbol{\phi}}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}+\mu_{d,2}{\mathbf{x}}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{x}}+\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1}+{\boldsymbol{\gamma}}^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (6)
where ${\boldsymbol{\gamma}}\in\mathbb{R}^{N(T+S+1)}$ is a Lagrange multiplier vector, and $\rho\in\mathbb{R}_{+}$ is a non-negative ADMM parameter. We solve (6) iteratively by minimizing variables ${\mathbf{x}}$ and ${\boldsymbol{\phi}}$ in alternating steps till convergence.
@@ 3.4.1 Minimizing ${\mathbf{x}}^{\tau+1}$
Fixing ${\boldsymbol{\phi}}^{\tau}$ and ${\boldsymbol{\gamma}}^{\tau}$ at iteration $\tau$, optimization in (6) for ${\mathbf{x}}^{\tau+1}$ becomes an unconstrained convex quadratic objective. The solution is a system of linear equations:
$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\mu_{u}{\mathbf{L}}^{u}+(\mu_{d,2}+\frac{\rho}{2}){\mathcal{L}}^{d}_{r}\right){\mathbf{x}}^{\tau+1}$ $=({\mathbf{L}}_{r}^{d})^{\top}(\frac{\rho}{2}\,{\boldsymbol{\phi}}^{\tau}+\frac{{\boldsymbol{\gamma}}^{\tau}}{2})+{\mathbf{H}}^{\top}{\mathbf{y}}.$ (7)
Because coefficient matrix ${\mathbf{H}}^{\top}{\mathbf{H}}+\mu_{u}{\mathbf{L}}^{u}+(\mu_{d,2}+\frac{\rho}{2}){\mathcal{L}}^{d}_{r}$ is sparse, symmetric and PD, (7) can be solved in linear time via conjugate gradient (CG) [29] without matrix inversion.
Splitting $\ell_{2}$-norm Terms: Instead of solving (7) directly for ${\mathbf{x}}^{\tau+1}$ at iteration $\tau$, we introduce again auxiliary variables ${\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in (5) for the $\ell_{2}$-norm GLR and DGLR terms respectively, and rewrite the optimization as
$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}{}{}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{z}}_{d}+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (8)
@@
$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{z}}_{d}+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (9)
$+({\boldsymbol{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|^{2}_{2}+({\boldsymbol{\gamma}}_{d}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{d})+\frac{\rho_{d}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{d}\|^{2}_{2}$
@@
where ${\boldsymbol{\gamma}}_{u},{\boldsymbol{\gamma}}_{d}\in\mathbb{R}^{N(T+S+1)}$ are multipliers, and $\rho_{u},\rho_{d}\in\mathbb{R}_{+}$ are ADMM parameters. Optimizing ${\mathbf{x}}$, ${\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in (9) in turn at iteration $\tau$ lead to three linear systems:
${\footnotesize\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right)}{\mathbf{x}}^{\tau+1}$ $={\small({\mathbf{L}}^{d}_{r})^{\top}(\frac{{\boldsymbol{\gamma}}^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}-\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau}+{\mathbf{H}}^{\top}{\mathbf{y}}}$ (10)
${\small\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right)}{\mathbf{z}}_{u}^{\tau+1}$ $={\small\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1}}$ (11)
@@
${\small\left(\mu_{d,2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}}{2}{\mathbf{I}}\right)}{\mathbf{z}}_{d}^{\tau+1}$ $={\small\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{x}}^{\tau+1}}.$ (12)
@@
Thus, instead of solving (7) for ${\mathbf{x}}^{\tau+1}$ directly at iteration $\tau$, ${\mathbf{x}}^{\tau+1}$, ${\mathbf{z}}_{u}^{\tau+1}$ and ${\mathbf{z}}_{d}^{\tau+1}$ are optimized in turn via (10) through (12) using CG. This splitting induces more network parameters for data-driven learning after algorithm unrolling (see Section 4.1) towards better performance, and affords us spectral filter interpretations of the derived linear systems.
Interpretation: To interpret solution ${\mathbf{z}}_{u}^{\tau+1}$ spectrally, we rewrite (11) as
@@
${\mathbf{z}}_{u}^{\tau+1}=\left(\frac{2\mu_{u}}{\rho_{u}}{\mathbf{L}}^{u}+{\mathbf{I}}\right)^{-1}\left(\frac{{\boldsymbol{\gamma}}_{u}}{\rho_{u}}+{\mathbf{x}}^{\tau+1}\right)={\mathbf{V}}\mathrm{diag}\left(\left\{\frac{1}{1+\frac{2\mu_{u}}{\rho_{u}}\lambda_{k}}\right\}\right){\mathbf{V}}^{\top}\left(\frac{{\boldsymbol{\gamma}}_{u}}{\rho_{u}}+{\mathbf{x}}^{\tau+1}\right)$ (13)
@@
where ${\mathbf{L}}^{u}={\mathbf{V}}\mathrm{diag}(\{\lambda_{k}\}){\mathbf{V}}^{\top}$ is the eigen-decomposition. Thus, using frequencies defined by eigen-pairs $\{(\lambda_{k},{\mathbf{v}}_{k})\}$ of ${\mathbf{L}}^{u}$, ${\mathbf{z}}_{u}^{\tau+1}$ is the low-pass filter output of ${\mathbf{x}}^{\tau+1}$ offset by $\frac{{\boldsymbol{\gamma}}_{u}}{\rho_{u}}$, with frequency response $f(\lambda)=(1+\frac{2\mu_{u}}{\rho_{u}}\lambda)^{-1}$.
As done for ${\mathbf{z}}_{u}^{\tau+1}$, we can rewrite ${\mathbf{z}}_{d}^{\tau+1}$ in (12) as
${\mathbf{z}}_{d}^{\tau+1}$ $={\mathbf{U}}\mathrm{diag}\left(\left\{\frac{1}{1+\frac{2\mu_{d,2}}{\rho_{d}}\xi_{k}}\right\}\right){\mathbf{U}}^{\top}\left(\frac{{\boldsymbol{\gamma}}_{d}}{\rho_{d}}+{\mathbf{x}}^{\tau+1}\right)$ (14)
@@
where ${\mathcal{L}}^{d}_{r}={\mathbf{U}}\mathrm{diag}(\{\xi_{k}\}){\mathbf{U}}^{\top}$ is the eigen-decomposition. Analogously, ${\mathbf{z}}_{d}^{\tau+1}$ is the low-pass filter output of ${\mathbf{x}}^{\tau+1}$ offset by $\frac{{\boldsymbol{\gamma}}_{d}}{\rho_{d}}$, with frequency response $f(\xi)=(1+\frac{2\mu_{d,2}}{\rho_{d}}\xi)^{-1}$.
To interpret solution ${\mathbf{x}}^{\tau+1}$ in (10) as a low-pass filter output, see Appendix B.

### v2

@@ 3.2.2 $\ell_{1}$-Norm Variational Term
Unlike DGLR in eq. 2, DGTV is not symmetric and has no obvious frequency interpretation. Nonetheless, we provide a two-channel filterbank interpretation in the sequel.
@@ 3.3 Optimization Formulation
where $\mu_{u},\mu_{d,2},\mu_{d,1}\in\mathbb{R}_{+}$ are weight parameters for the three regularization terms. Combination of $\ell_{2}$- and $\ell_{1}$-norm penalties—DGLR and DGTV for directed graph ${\mathcal{G}}^{d}$ in our case [^5] We can employ both $\ell_{2}$- and $\ell_{1}$-norm regularization terms for the undirected graph ${\mathcal{G}}^{u}$ as well. For simplicity, we focus on designing new regularization terms for directed graphs in this paper. —is called elastic net regularization in statistics (Zou & Hastie, 2005) with demonstrable improved robustness. Further, employing both regularization terms means that, after algorithm unrolling, we can adapt an appropriate mixture by learning weights $\mu_{d,2}$ and $\mu_{d,1}$ per neural layer.
Convex minimization in eq. 4 is composed of smooth $\ell_{2}$-norm terms and one non-smooth $\ell_{1}$-norm term. We pursue a divide-and-conquer approach and solve eq. 4 via an ADMM framework (Boyd et al., 2011) by minimizing $\ell_{2}$- and $\ell_{1}$-norm terms in eq. 4 alternately.
@@ 3.4 ADMM Optimization Algorithm
We first introduce auxiliary variable $\boldsymbol{\phi}$} and rewrite the optimization in eq. 4 as
@@
Using the augmented Lagrangian method (Boyd & Vandenberghe, 2004), we rewrite eq. 5 in an unconstrained form:
$\min_{{\mathbf{x}},{\boldsymbol{\phi}}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{x}}^{\top}{\mathbf{L}}^{u}{\mathbf{x}}+\mu_{d,2}{\mathbf{x}}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{x}}+\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1}+{\bm{\gamma}}^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (6)
where ${\bm{\gamma}}\in\mathbb{R}^{N(T+S+1)}$ is a Lagrange multiplier vector, and $\rho\in\mathbb{R}_{+}$ is a non-negative ADMM parameter. We solve eq. 6 iteratively by minimizing $\mathbf{x}$} and $\boldsymbol{\phi}$} in alternating steps till convergence.
@@ 3.4.1 Minimizing ${\mathbf{x}}^{\tau+1}$
Fixing ${\boldsymbol{\phi}}^{\tau}$ and ${\bm{\gamma}}^{\tau}$ at iteration $\tau$, optimization in eq. 6 for ${\mathbf{x}}^{\tau+1}$ becomes an unconstrained convex quadratic objective. The solution is a system of linear equations:
$\hskip-10.00002pt\left({\mathbf{H}}^{\top}{\mathbf{H}}+\mu_{u}{\mathbf{L}}^{u}+(\mu_{d,2}+\frac{\rho}{2}){\mathcal{L}}^{d}_{r}\right){\mathbf{x}}^{\tau+1}$ $=({\mathbf{L}}_{r}^{d})^{\top}(\frac{\rho}{2}\,{\boldsymbol{\phi}}^{\tau}+\frac{{\bm{\gamma}}^{\tau}}{2})+{\mathbf{H}}^{\top}{\mathbf{y}}.$ (7)
Because coefficient matrix ${\mathbf{H}}^{\top}{\mathbf{H}}+\mu_{u}{\mathbf{L}}^{u}+(\mu_{d,2}+\frac{\rho}{2}){\mathcal{L}}^{d}_{r}$ is sparse, symmetric and PD, eq. 7 can be solved in linear time via conjugate gradient (CG) (Shewchuk, 1994) without matrix inversion.
Splitting $\ell_{2}$-norm Terms: Instead of solving eq. 7 directly for ${\mathbf{x}}^{\tau+1}$ at iteration $\tau$, we introduce again auxiliary variables ${\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in eq. 5 for the $\ell_{2}$-norm GLR and DGLR terms respectively, and rewrite the optimization as
$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{z}}_{d}+({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (8)
@@
$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|^{2}_{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}_{r}^{d}{\mathbf{z}}_{d}+({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|^{2}_{2}$ (9)
$+({\bm{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|^{2}_{2}+({\bm{\gamma}}_{d}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{d})+\frac{\rho_{d}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{d}\|^{2}_{2}$
@@
where ${\bm{\gamma}}_{u},{\bm{\gamma}}_{d}\in\mathbb{R}^{N(T+S+1)}$ are multipliers, and $\rho_{u},\rho_{d}\in\mathbb{R}_{+}$ are ADMM parameters. Optimizing $\mathbf{x}$}, ${\mathbf{z}}_{u}$ and ${\mathbf{z}}_{d}$ in eq. 9 in turn at iteration $\tau$ lead to three linear systems:
@@
$\hskip-4.2679pt{\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right)}{\mathbf{x}}^{\tau+1}$ $={({\mathbf{L}}^{d}_{r})^{\top}\!\left(\frac{{\bm{\gamma}}^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau}\right)-\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}-\frac{{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau}+{\mathbf{H}}^{\top}{\mathbf{y}}}$ (10)
${\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right)}{\mathbf{z}}_{u}^{\tau+1}$ $={\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1}}$ (11)
${\left(\mu_{d,2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}}{2}{\mathbf{I}}\right)}{\mathbf{z}}_{d}^{\tau+1}$ $={\frac{{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{x}}^{\tau+1}}.$ (12)
@@
Thus, instead of solving eq. 7 for ${\mathbf{x}}^{\tau+1}$ directly at iteration $\tau$, ${\mathbf{x}}^{\tau+1}$, ${\mathbf{z}}_{u}^{\tau+1}$ and ${\mathbf{z}}_{d}^{\tau+1}$ are optimized in turn via eq. 10 through eq. 12 using CG. This splitting induces more network parameters for data-driven learning after algorithm unrolling (see Section 4.1) towards better performance, and affords us spectral filter interpretations of the derived linear systems.
Interpretation: To interpret solution ${\mathbf{z}}_{u}^{\tau+1}$ spectrally, we rewrite eq. 11 as
@@
$\hskip-2.84526pt{\mathbf{z}}_{u}^{\tau+1}=\left(\frac{2\mu_{u}}{\rho_{u}}{\mathbf{L}}^{u}+{\mathbf{I}}\right)^{-1}\left(\frac{{\bm{\gamma}}_{u}}{\rho_{u}}+{\mathbf{x}}^{\tau+1}\right)={\mathbf{V}}\mathrm{diag}\left(\left\{\frac{1}{1+\frac{2\mu_{u}}{\rho_{u}}\lambda_{k}}\right\}\right){\mathbf{V}}^{\top}\left(\frac{{\bm{\gamma}}_{u}}{\rho_{u}}+{\mathbf{x}}^{\tau+1}\right)$ (13)
@@
where ${\mathbf{L}}^{u}={\mathbf{V}}\mathrm{diag}(\{\lambda_{k}\}){\mathbf{V}}^{\top}$ is the eigen-decomposition. Thus, using frequencies defined by eigen-pairs $\{(\lambda_{k},{\mathbf{v}}_{k})\}$ of ${\mathbf{L}}^{u}$, ${\mathbf{z}}_{u}^{\tau+1}$ is the low-pass filter output of ${\mathbf{x}}^{\tau+1}$ offset by $\frac{{\bm{\gamma}}_{u}}{\rho_{u}}$, with frequency response $f(\lambda)=(1+\frac{2\mu_{u}}{\rho_{u}}\lambda)^{-1}$.
As done for ${\mathbf{z}}_{u}^{\tau+1}$, we can rewrite ${\mathbf{z}}_{d}^{\tau+1}$ in eq. 12 as
${\mathbf{z}}_{d}^{\tau+1}$ $={\mathbf{U}}\mathrm{diag}\left(\left\{\frac{1}{1+\frac{2\mu_{d,2}}{\rho_{d}}\xi_{k}}\right\}\right){\mathbf{U}}^{\top}\left(\frac{{\bm{\gamma}}_{d}}{\rho_{d}}+{\mathbf{x}}^{\tau+1}\right)$ (14)
@@
where ${\mathcal{L}}^{d}_{r}={\mathbf{U}}\mathrm{diag}(\{\xi_{k}\}){\mathbf{U}}^{\top}$ is the eigen-decomposition. Analogously, ${\mathbf{z}}_{d}^{\tau+1}$ is the low-pass filter output of ${\mathbf{x}}^{\tau+1}$ offset by $\frac{{\bm{\gamma}}_{d}}{\rho_{d}}$, with frequency response $f(\xi)=(1+\frac{2\mu_{d,2}}{\rho_{d}}\xi)^{-1}$.
To interpret solution ${\mathbf{x}}^{\tau+1}$ in eq. 10 as a low-pass filter output, see Appendix C.